OpenStudy (anonymous):

Solve each problem. Please round to the nearest tenth if you need to. *medal for whoever answers* 36. What is the radius of a circle that has a circumference of 37 m? 37. what is the width of a rectangle that has a length of 10m and an area of 65m^2? 38. a rectangle has a perimeter of 20cm and a length of 6 cm. what is the width?

OpenStudy (anonymous):

"36. What is the radius of a circle that has a circumference of 37 m? \[37 = 2*\pi*r\]solve for r "37. what is the width of a rectangle that has a length of 10m and an area of 65m^2? \[65 = 10*w\]solve for w "38. a rectangle has a perimeter of 20cm and a length of 6 cm. what is the width? \[20 = 2*6 + 2*w\] solve for w

OpenStudy (anonymous):

thanku

OpenStudy (anonymous):

love the name...prettyfeather...lol

OpenStudy (anonymous):

oh...it didn't let me type it in